OpenStudy (anonymous):

sina_____=1/2(sin(a+b)+sin(a-b)).................is which of these......cosb, sinb, sina, cosa

OpenStudy (anonymous):

cosb

OpenStudy (anonymous):

Using sum and difference identities to expand the values in the 1/2: \[(\sin a \cos b + \cos a \sin b)+(\sin a \cos b - \cos a \sin b)\] Simplify: \[2\sin a\cos b\] Put it back in the equation: \[\frac{1}{2}(2\sin a \cos b)\] 2's cancel: \[\sin a \cos b\] Therefore, \(\cos b\)!

OpenStudy (anonymous):

thank u so much!